OpenStudy (anonymous):

Help please? Simple equation :) 2y+1.8=4y-4.4

OpenStudy (xxdragoonxx):

2y+1.8=4y-4.4 1.8=4y-4.4-2y = 1.8=2y-4.4 = 1.8+4.4=2y = 6.2=2y = 6.2/2=y 3.1=y y=3.1